qwevery:

August hosted two dinner parties for his friends. Twenty-five guests attended the first party, and thirty-three guests attended the second party. What is the percentage increase of the number of guests from the first party to the second party? A. 80% B. 68% C. 32% D. 20%

mdobbs6856:

Increase % = Increase / Original Number *100 Increase % = 6/25 * 100 Increase % = 6 * 4 Increase % =24

qwevery:

lol thank u:)

mdobbs6856:

It would round down to 20%

mdobbs6856:

YW!

mdobbs6856:

Oh! Also, Welcome to QuestionCove!

qwevery:

thank u:)

mdobbs6856:

@qwevery wrote:
thank u:)
Np